OpenStudy (nuccioreggie):

Betsey is 11 years older than Waldo. Betsey's age is 15 years less than three times Waldo's age. The system below models the relationship between Betsey's age (b) and Waldo's age (w): b = w + 11 b = 3w − 15 Which of the following methods is correct to find Betsey's and Waldo's age? Solve w + 11 = 3w − 15 to find the value of w. Solve b + 11 = 3b − 15 to find the value of b. Write the points where the graphs of the equations intersect the x-axis. Write the points where the graphs of the equations intersect the y-axis.

OpenStudy (nuccioreggie):

@whovianchick

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (anonymous):

Lets solve. w+11=3w-15 2w=26 w=13 b=w+11 b=(13)+11=24

OpenStudy (anonymous):

Your answer is 24.

OpenStudy (nuccioreggie):

so which is the answer choice

OpenStudy (anonymous):

What do you think?

OpenStudy (nuccioreggie):

A

OpenStudy (michele_laino):

from the first equation we get the value of \(b\), which is \(b=w+11\). Please replace \(b\) in the second equation with such value

OpenStudy (mathmale):

My suggestion is that you choose a method (such as substitution) of solving this system of linear equations, and then proceed with it. Then, compare your efforts to those listed in the four answer choices. That may help you choose the correct choice. Ladies and Gentlemen: PLEASE stop giving out direct answers without the involvement of the person who asked the question!! Provide guidance, not answers.

OpenStudy (nuccioreggie):

THE ANSWER IS A